OpenStudy (anonymous):

5.Find the simplified form of the expression. Give your answer in scientific notation. (3 x 10^-6) (6 x 10^-8)

OpenStudy (anonymous):

use a calculator.. that just a multiplication of two numbers... 1*10^-2 = 0.01 and 1*10^2 = 100

OpenStudy (campbell_st):

this is multiplying 1 numbers written in scientific notation so its \[1.8 \times 10 ^ {-13}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you know campbell.. you really should give answers to questions that easy... explaining how to do it, okay; but this is supposed to be a study site, not a let's be lazy and just ask the answers to all my homework questions :P how will we ever learn if we never have to understand... (A)